OpenStudy (killercboy):

Mike is driving a total of 100 miles. He drove 23 miles before stopping at a gas station. Then he drove 58 more miles before stopping for lunch. About how much farther does Mike have to drive?

OpenStudy (shamim):

58+23=?

OpenStudy (shamim):

58+23=81?

OpenStudy (shamim):

100-81=?

OpenStudy (shamim):

Response plz!!!!!!!!